OpenStudy (anonymous):

How do i find the angle in degree measure. I have a few problems like this..How do i solve them? Example of one of them is 27pi/10

jimthompson5910 (jim_thompson5910):

Multiply the given radian angle by \[\large \frac{180}{\pi}\] to get \[\large \frac{27\pi}{10}\times\frac{180}{\pi}=486\] So the angle in degrees is 486 degrees

OpenStudy (anonymous):

Thanks!. Question: Would the answer to 11pi/8 be 247.5 degrees?

jimthompson5910 (jim_thompson5910):

yep you got it

OpenStudy (anonymous):

What about converting degrees to radian measure.. I have 12 degrees.. Would it end up as 1/15 or what?

jimthompson5910 (jim_thompson5910):

Flip the fraction around and instead multiply by \[\large \frac{\pi}{180}\]